What is the standard form with the given zeros 5, -3, -2




Answer by solve_for_x(190)

Since the zeros are 5, -3 and -2, the factors of the polynomial are:

(x - 5), (x + 3), and (x + 2)

Setting the function equal to the product of the factors gives you:

f(x) = (x - 5)(x + 3)(x + 2)

f(x) = (x - 5)(x^2 + 5x + 6)

f(x) = x^3 + 5x^2 + 6x - 5x^2 - 25x - 30

f(x) = x^3 - 19x - 30